write the equation of the line that passes through (10,-4) and is perpendicular to y=5x + 3 in slope intercept form
Answer by stanbon(75887)

write the equation of the line that passes through (10,-4) and is perpendicular to y=5x + 3 in slope intercept form
--------
The given line has slope = 5.
---
Your equation must have slope = -1/5.
-----
Form: y = mx + b
-4 = (-1/5)10 + b
-4 = -2 + b
b = -2
--------
Equation:
y = (-1/5)x -2
=================
Cheers,
Stan H.
=======================